OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. x + 8y = 14 -3x + 4y = 0

OpenStudy (mertsj):

Can you multiply the first equation by 3?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

x + 8y = 14 x + 8 *(+ 0 + 0.75x )=+ 14 x + 0 + 6x =+ 14 7x + 0 =+ 14 7x + 0 =+ 14 - 0 - 0 7x =+ 14 ------ ----- + 7 + 7 x = 2

OpenStudy (mertsj):

I'm pretty sure you used substitution and not elimination as the directions said.

OpenStudy (anonymous):

i think i did still trying to get used to the elimination method.

OpenStudy (mertsj):

Well then, why didn't you do what i said and multiply the first equation by 3?

OpenStudy (anonymous):

my newc answers i came out with was x=7/3 and y=7/4

OpenStudy (zepp):

http://puu.sh/pQbD

OpenStudy (zepp):

x should be 2 and y should be 1.5 (or 1/2)

OpenStudy (mertsj):